Solve for y: (1/3)(3y + 12) = – 5y – 4(y – 6)



Solve for x: (4 + 3x)/5 + (7 – x)/3 = (7/15)

Answer by jai_kos(139)

Given (1/3)(3y + 12) = – 5y – 4(y – 6)
y + 4 = -5y -4y + 24
y + 4 = - 9y + 24
y + 9y = 24 - 4
10 y = 20
y = 20 /10
y = 2

Solve for x: (4 + 3x)/5 + (7 – x)/3 = (7/15)
Given (4 + 3x)/5 + (7 – x)/3 = (7/15)
Multiply the above equation with 15, we get
(4 + 3x) * 3 + ( 7 - x) * 5 = 7
12 + 9x + 35 - 5x = 7
47 + 4x -7 = 0
40 + 4x = 0
10 + x = 0
x = -10